OpenStudy (anonymous):

Find the slope-intercept form of the equation of the line passing through the points. (3, 2), (−3, −3) i got 1/2???

OpenStudy (anonymous):

slope = (-3 - 2)/(-3 - 3) = 5/6

OpenStudy (anonymous):

it says thats wrong :(

OpenStudy (karatechopper):

lets check

OpenStudy (karatechopper):

can u show me how u did tht mlhuyck

OpenStudy (karatechopper):

because tht is wrong the answer u came up with..

OpenStudy (anonymous):

i got y=5/6x +1/2

OpenStudy (karatechopper):

tht is correct^

OpenStudy (anonymous):

you have to find slope interecept form not the slope first you find the slope which was indeed 5/6 then you you use point slope form to fine slope intercept form y-y1=m(x-x1) so y-2=5/6(x-3) thats how i got y=5/6x+1/2

OpenStudy (anonymous):

it says thats wrong :(

OpenStudy (anonymous):

Thank you thank you!!